OpenStudy (anonymous):

When factoring a quadratic equation in standard form ax²+bx+c = 0 and a is not greater than 1, the b and c share a special relationship. To all; Share with the special relationship b and c share.

OpenStudy (dumbcow):

If it factors to (x+p)(x+q) then b = p+q c = p*q

myininaya (myininaya):

whats the question?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

(-2pb)^2/4p +k

OpenStudy (anonymous):

What is the special relationship b & c share?

myininaya (myininaya):

so a can be less than one but not greater?

OpenStudy (anonymous):

to my understanding that is correct myininaya

myininaya (myininaya):

hey i can attach something pretty cool about this if you give me a second ok?

myininaya (myininaya):

like i found a new way to prove the quadratic formula and it might help you see the relation between a, b, and c

myininaya (myininaya):

you just really need the first part ok?

myininaya (myininaya):

but you can read the whole thing if you like

OpenStudy (anonymous):

myininaya, this is sorta helpful, ia ma extremely math (retarded) and although I see lots that seems like it answers my question. can you put it "stupid" terms for me? all i need is the special relationship...

OpenStudy (dumbcow):

thats pretty good myininaya, i think i prefer the completing the square proof though :)

myininaya (myininaya):

lol okay lets see stupid terms: what you want to do is find factors of a*c that add up to be b

myininaya (myininaya):

find two factors of a*c that have product a*c and have sum b*

myininaya (myininaya):

thanks cow :)

OpenStudy (anonymous):

factoring a quadratic equation in standard form ax²+bx+c = 0 and a is not greater than 1, the b and c share a special relationship. so in order to find the "special relationship" you have to solve it first?

myininaya (myininaya):

i think your question means if a=1 i see that it says a is not greater than 1 which means it could be less than 1 or equal to? it doesn't make sense for it to exclude positive a and not negative a

myininaya (myininaya):

\[x^2+bx+c\] say the roots are \[r_1 \& r_2\] \[x^2+bx+c=(x-r_1)(x-r_2)=x^2-r_2x-r_1x+r_1r_2=x^2-(r_1+r_1)x+r_1r_2\]

OpenStudy (anonymous):

i guess another question is: is there a standard "special relationship" between b & c in a quadratic equation OR does it vary by equation? perhaps i am overthinking this?

myininaya (myininaya):

\[b=r_1+r_2 \& c=r_1r_2\]

myininaya (myininaya):

and this is for a=1 okay?

myininaya (myininaya):

i think what cow has is good

OpenStudy (dumbcow):

is there a way to put b in terms of c...is that what you are looking for

OpenStudy (dumbcow):

b equals the sum of factors of c...theres the special relationship (a=1)

myininaya (myininaya):

\[ax^2+bx+c=a(x^2+\frac{b}{a}x+\frac{c}{a})\] so we have roots \[r_1 \& r_2\] => \[a(x^2+\frac{b}{a}x+\frac{c}{a})=a(x-r_1)(x-r_2)=a(x^2+-(r_1+r_2)+r_1r_2)\] => \[\frac{b}{a}=-r_1+r_2 \& \frac{c}{a}=r_1r_2\]

myininaya (myininaya):

oops there should be paranthesis around (r_1+r_2)

myininaya (myininaya):

\[\frac{b}{a}=-(r_1+r_2) \& \frac{c}{a}=r_1r_2\]

myininaya (myininaya):

there that looks better

OpenStudy (anonymous):

=) i really appreciate your assistance!